Please help with this:
Solve for x: (x+4)^2-3=6
Can you please show me how you got the answer?
Thank You

Answer by jim_thompson5910(35256)

(x+4)^2-3=6

(x+4)^2-3+3=6+3 ... Add 3 to both sides.

(x+4)^2 = 9

sqrt( (x+4)^2 ) = +-sqrt( 9 ) ... Take the square root of both sides

x + 4 = 3 or x + 4 = -3

x + 4 - 4 = 3-4 or x + 4-4 = -3-4

x = -1 or x = -7